# WiMANS: A Benchmark Dataset for WiFi-based Multi-user Activity Sensing

### 24 Jan 2024 v1

@@ 5 Discussion
To the best of our knowledge, WiMANS is the first WiFi-based multi-user activity sensing dataset, containing the CSI of dual WiFi bands with synchronized videos from various environments. Despite its great promise for WiFi-based human sensing, we acknowledge some limitations. Limitations. (1) Daily Activities: WiMANS includes common daily activities only, not ethically risky activities (e.g., falling, fighting), because we regard WiMANS as a generic multi-user sensing dataset, and the current activities in WiMANS are representative of daily life. We will consider other activities in our future work, given ethical approval to ensure the safety of volunteers. (2) More Challenging Conditions: WiFi-based human sensing has the potential to solve the issues of obstacles. However, we do not intentionally set up obstacles since we also use a camera to capture videos, though WiMANS does contain occlusion due to obstacles (e.g., chairs, desks) and indeed people also overlap with each other. In our future work, we will further analyze the impacts of obstacles on WiFi-based multi-user sensing. Future Work. (1) Dual-band Augmented Sensing: 2.4 GHz and 5 GHz WiFi bands have different advantages for human sensing, as mentioned in 3.2. Hence, we plan to investigate multi-user sensing augmented by dual WiFi bands for better performance. (2) Cross-modal Supervision: Given that WiFi-based models still have vast room for improvement compared with video-based models, we plan to utilize video data to improve WiFi-based models though cross-modal supervision. (3) Pose Estimation: Since WiMANS collects both WiFi CSI and videos, we plan to extend our dataset by annotating videos with human body joints and perform WiFi-based multi-user pose estimation. Ethics Statement. Human subject study in WiMANS has been reviewed and approved by institutional ethics committee (IRB). Prior to data collection, each subject has been given the details of WiMANS and signed the consent form regarding safety, privacy, and releasing identifiable information. The activities in WiMANS are common in daily life (e.g., Walking) to minimize participant risks. We create aliases for all subjects (e.g., User 1) and blur their faces in videos to anonymize personally identifiable information. WiMANS should be used for academic research purposes only. We provide further discussions in Appendix D.
@@ 6 Conclusion
In this paper, we introduce WiMANS, the first WiFi-based human sensing dataset that incorporates multiple users in each CSI sample. WiMANS contains 11286 CSI samples of dual WiFi bands and synchronized videos, annotated with user identities, locations, and activities. We run benchmark experiments on WiMANS to evaluate the multi-user sensing performance of existing models based on WiFi and videos. WiFi-based models show desirable performance for multi-user identification and localization, while there is still vast room for improvement in activity recognition compared with video-based models. We make all data, source code, and documents publicly available, anticipating that WiMANS can advance WiFi-based human sensing research from single-user sensing towards more practical multi-user sensing. We hope that WiMANS can enable a variety of future works, such as dual-band augmented sensing, cross-modal supervision, and multi-user pose estimation.
Supplementary Material
Appendix A Data Collection
Herein, we provide the details of data collection in WiMANS, including the description of activities (A.1), user statistics (A.2), and the details of 114 user groups (A.3).
A.1 Description of Activities
WiMANS consists of 9 daily activities: (a) Nothing, (b) Walking, (c) Rotation, (d) Jumping, (e) Waving, (f) Lying Down, (g) Picking Up, (h) Sitting Down, (i) Standing Up. Supplement Figure 7 presents the descriptions and examples of these activities.
@@ A.1 Description of Activities
Figure 7: Descriptions and examples of 9 daily activities in WiMANS.
A.2 User Statistics
@@
Table 5: Statistics of 6 users in WiMANS.
A.3 User Groups
In WiMANS, we organize volunteers into user groups which correspond to different numbers of user and environments. In each group, each user holds different scripts (i.e., $\bm{\alpha}$, $\bm{\beta}$, $\bm{\delta}$, $\bm{\gamma}$, and $\bm{\lambda}$) to perform activities independently yet simultaneously. Supplement Tables 6 $\sim$ 12 present the details of 114 user groups in WiMANS.
@@ 18 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Table 6: User group 1$\sim$18 in the classroom using the 2.4 GHz WiFi band.
@@ 36 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Table 7: User group 19$\sim$36 in the classroom using the 5 GHz WiFi band.
@@ 54 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Table 8: User group 37$\sim$54 in the meeting room using the 2.4 GHz WiFi band.
@@ 72 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Table 9: User group 55$\sim$72 in the meeting room using the 5 GHz WiFi band.
@@ 90 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Table 10: User group 73$\sim$90 in the empty room using the 2.4 GHz WiFi band.
@@ 99 Empty Room $\bm{\lambda}$ $\bm{\alpha}$
Table 11: User group 91$\sim$108 in the empty room using the 5 GHz WiFi band.
@@
Table 12: User group 201$\sim$206 with 0 user.
Appendix B Data Accessibility, Maintenance, and Usage
@@
  4. Raw CSI data are saved in “*.mat” files, which can be read using SciPy.
    The preprocessed data of CSI amplitude are saved in “*.npy” files, which can be read using NumPy.
    Video data are saved in “*.mp4” files, which can be read using PyTorch.
    Annotations are saved in the “annotation.csv” file, which can be read using Pandas.
  5. Details of how to use the dataset and source code are provided in the README.md file.
  6. We will provide and maintain all data, source code, and documents on the public project page.
Appendix C Experiments

### 12 Mar 2024 v2

@@ 5 Discussion
To the best of our knowledge, WiMANS is the first benchmark dataset for WiFi-based multi-user activity sensing. In this section, we discuss the limitations and future work of WiMANS and provide an ethics statement.
5.1 Limitations
WiMANS aims to represent the most common multi-user scenarios in daily life. Therefore, we discuss the most general settings but not specific activities, conditions, dedicated devices, etc. Such general settings may lead to some limitations. Daily Activities. WiMANS includes common daily activities only, not ethically risky activities (e.g., falling, fighting), and the current activities in WiMANS are representative of daily life [73, 77, 3]. We will further consider other activities in our future work, given ethical approval to ensure the safety of volunteers. Challenging Conditions. WiFi-based human sensing has the potential to solve the issues of obstacles. However, we do not intentionally set up obstacles since we use a camera to capture videos, though WiMANS does contain occlusion due to obstacles (e.g., chairs, desks) and indeed people also overlap with each other. WiFi Devices. We collect WiFi CSI using a single transmitter-receiver pair equipped with Intel 5300 Network Interface Cards, the most commonly used devices in the previous works [73, 39, 44, 22, 77, 3]. More transmitter-receiver pairs may benefit WiFi-based multi-user sensing [1], while other dedicated devices can also collect CSI using the Atheros CSI tool [70, 65], the AX-CSI tool [40, 20], etc.
5.2 Future Work
Multi-user Pose Estimation. Recent works have discussed WiFi-based pose estimation [80, 79, 35, 60, 30, 18, 47], but they have not released their datasets. Since WiMANS collects both WiFi CSI and videos, we can extend our dataset by annotating videos with human body joints for multi-user pose estimation. Dual-band Augmented Sensing. 2.4 GHz and 5 GHz WiFi bands have different advantages for human sensing, as mentioned in 3.2. We plan to investigate multi-user sensing augmented by dual WiFi bands for better performance. Cross-domain Sensing. Recent studies [6, 83] have discussed cross-domain single-user sensing based on WiFi CSI. We plan to explore cross-domain multi-user sensing with varying numbers of users, unseen environments/users, etc.
5.3 Ethics Statement
Human subject study in WiMANS has been reviewed and approved by institutional ethics committee (IRB). Prior to data collection, each subject has been given the details of WiMANS and signed the consent form regarding safety, privacy, and releasing identifiable information. The activities in WiMANS are common in daily life (e.g., Walking) to minimize participant risks. We create aliases for all subjects (e.g., User 1) and blur their faces in videos to anonymize personally identifiable information. WiMANS should be used for academic research purposes only. We provide further discussions in Appendix D.
@@ 6 Conclusion
In this paper, we introduce WiMANS, the first WiFi-based human sensing dataset that incorporates multiple users in each CSI sample. WiMANS contains 11286 CSI samples of dual WiFi bands and synchronized videos, annotated with user identities, locations, and activities. We run benchmark experiments on WiMANS to evaluate the multi-user sensing performance of existing models based on WiFi and videos. WiFi-based models show desirable performance for multi-user identification and localization, while there is still vast room for improvement in activity recognition compared with video-based models. We make all data, source code, and documents publicly available, anticipating that WiMANS can advance WiFi-based human sensing research from single-user sensing towards more practical multi-user sensing. We hope that WiMANS can enable a variety of future works, such as multi-user pose estimation, dual-band augmented sensing, and cross-domain sensing.
Appendix
Appendix 0.A Data Collection
Herein, we provide more details of data collection in WiMANS, including the description of activities (0.A.1), user statistics (0.A.2), and the details of 114 user groups (0.A.3).
0.A.1 Description of Activities
WiMANS consists of 9 daily activities: (a) Nothing, (b) Walking, (c) Rotation, (d) Jumping, (e) Waving, (f) Lying Down, (g) Picking Up, (h) Sitting Down, (i) Standing Up. Supplement Fig. 1 presents the descriptions and examples of these activities.
@@ 0.A.1 Description of Activities
Supplement Fig. 1: Descriptions and examples of 9 daily activities in WiMANS.
0.A.2 User Statistics
@@
0.A.3 User Groups
In WiMANS, we organize volunteers into user groups which correspond to different numbers of user and environments. In each group, each user holds different scripts (i.e., $\bm{\alpha}$, $\bm{\beta}$, $\bm{\delta}$, $\bm{\gamma}$, and $\bm{\lambda}$) to perform activities independently yet simultaneously. Supplement Tables 2 $\sim$ 8 present the details of 114 user groups in WiMANS.
Supplement Table 2: User group 1$\sim$18 in the classroom using the 2.4 GHz WiFi band.
@@ 18 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Supplement Table 3: User group 19$\sim$36 in the classroom using the 5 GHz WiFi band.
@@ 36 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Supplement Table 4: User group 37$\sim$54 in the meeting room using the 2.4 GHz WiFi band.
@@ 54 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Supplement Table 5: User group 55$\sim$72 in the meeting room using the 5 GHz WiFi band.
@@ 72 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Supplement Table 6: User group 73$\sim$90 in the empty room using the 2.4 GHz WiFi band.
@@ 90 $\bm{\alpha}$ $\bm{\delta}$ $\bm{\gamma}$ $\bm{\lambda}$ $\bm{\beta}$
Supplement Table 7: User group 91$\sim$108 in the empty room using the 5 GHz WiFi band.
@@ 99 Empty Room $\bm{\lambda}$ $\bm{\alpha}$
Supplement Table 8: User group 201$\sim$206 with 0 user.
@@
Appendix 0.B Data Accessibility, Maintenance, and Usage
@@
The preprocessed data of CSI amplitude are saved in “*.npy” files, which can be read using NumPy. Video data are saved in “*.mp4” files, which can be read using PyTorch. Annotations are saved in the “annotation.csv” file, which can be read using Pandas. Details of how to use the dataset and source code are provided in the README.md file. We will provide and maintain all data, source code, and documents on the public project page.
Appendix 0.C Experiments
@@ 0.C.1 Implementation Details
Supplement Table 9: Hyperparameters of WiFi-based and video-base models.